find (f*g)(-4) and (g*f)(-4)
f(x)=-2x-1; g(x)=(x^2)+5

Answer by Edwin McCravy(20056)

I fear that you are typing an asterisk * when you should be 
typing a small round circle ◦ which means "composed with",
referring to the so-called "composition" of two functions.

f◦g(x) and f*g(x) are entirely different things.  f*g(x)
means to multiply the expressions together, but f◦g(x)
means to substitute what g(x) equals for x in what f(x) equals.

To get the little circle ◦ on here, type this &# 9702;  
but don't skip a space between the # and the 9.  I skipped a
space above so you could see what characters I typed insted
of just another ◦. 

find (f◦g)(-4) and (g◦f)(-4)
f(x)=-2x-1; g(x)=x²+5

f◦g(-4) means f(g(-4))

So first find g(-4):

g(-4) = (-4)²+5 = 16+5 = 21

So

f◦g(-4) = f(g(-4)) = f(21) = -2(21)-1 = -42-1 = -43

---------

g◦f(-4) means g(f(-4))

So first find f(-4):

f(-4) = -2(-4)-1 = 8-1 = 7

So

g◦f(-4) = g(f(-4)) = g(7) = (7)²+5 = 49+5 = 54


Edwin